OpenStudy (mathmath333):

how do i find the units digit of \(\LARGE \begin{align} \color{black}{ 58^{{52}^{98}} \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (misty1212):

it is either 2, 4, 6, or 8 right?

OpenStudy (mathmath333):

YES

OpenStudy (misty1212):

pattern would look like \[8,4,2,6\]

OpenStudy (misty1212):

since 52 is divisible by 4, the units place in \(8^{52}\) is \(6\)

OpenStudy (misty1212):

\[\overbrace{8}^1\overbrace{4}^2\overbrace{2}^3\overbrace{6}^4\]

OpenStudy (misty1212):

repeat the process for \(6^{96}\)

OpenStudy (mathmath333):

how did \(\LARGE 6^{96}\) came from

OpenStudy (misty1212):

i guess a mistake on my part , should be \[6^{98}\]

OpenStudy (mathmath333):

YES how did that came

OpenStudy (misty1212):

because the units digit of \(8^{52}\) is \(6\)

OpenStudy (misty1212):

is that much clear, or is that confusing?

ganeshie8 (ganeshie8):

\[58^{{52}^{98}} = (10x-2)^{4y} \equiv (-2)^{4y}\equiv 6^y \pmod{10}\]

OpenStudy (mathmath333):

yes it was clear thnx

OpenStudy (mathmath333):

that look familier to binomial :D

ganeshie8 (ganeshie8):

yes, to simplify matter we can replace 10 by 0 in mod 10

ganeshie8 (ganeshie8):

\[58^{{52}^{98}} = (6\times10 - 2)^{{52}^{98}} \equiv (6\times 0-2)^{{52}^{98}} = (-2)^{{52}^{98}} \pmod{10}\]

ganeshie8 (ganeshie8):

noticing 52 is divisible by 4 allows you to conclude 52^98 is divisible by 4 so we have 52^98 = 4y : \[ (-2)^{{52}^{98}} = (-2)^{4y}\]

OpenStudy (mathmath333):

@misty1212 ,how did i find \(\huge 2^{{117}^{188}}\) do i need to find the units digit of \(\huge 2^{117}\) first

ganeshie8 (ganeshie8):

@satellite73

ganeshie8 (ganeshie8):

either of them should be able to assist you :)

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{ \dfrac{2^{117}}{10} \implies 2\hspace{.33em}\\~\\ and \hspace{.33em}\\~\\ \dfrac{2^{188}}{10} \hspace{.33em}\\~\\ \implies 6\hspace{.33em}\\~\\ }\end{align}\) but the answer is 2

OpenStudy (mathmath333):

ok this method should work \(\LARGE \begin{align} \color{black}{ 2^{{117}^{188}} \hspace{.33em}\\~\\ \implies 2^{\frac{117^{188}}{4}} \hspace{.33em}\\~\\ \implies 2^{\frac{1^{188}}{4}} \hspace{.33em}\\~\\ \implies 2^{\frac{1}{4}} \hspace{.33em}\\~\\ \implies 2^{1} \hspace{.33em}\\~\\ \implies 2 \hspace{.33em}\\~\\ }\end{align}\)

ganeshie8 (ganeshie8):

we may write \(117\) as \(4k+1\) and reduce

ganeshie8 (ganeshie8):

\[\large 2^{{117}^{188}} =2^{(4k+1)^{188}} \]

OpenStudy (mathmath333):

I am applying strategy of \((mod~~4)\) and them finding the equivalent digit to \(\large 2^{mod}\)

ganeshie8 (ganeshie8):

\[\large 2^{{117}^{188}} =2^{(4k+1)^{188}} = 2^{4m+1} = 2*2^{4m} \]

OpenStudy (mathmath333):

yes that \(6^{n}\) trick is powerful , but various method confuse me

ganeshie8 (ganeshie8):

there is a mistake in your method, let me show it : how do you work the units digit of \(3^{117^{188}}\) using your method ?

OpenStudy (mathmath333):

first i did \( 117^{188} ~~~~(mod ~~4)\) which was \( 1\) \(so\\~\\ 2^1\)

ganeshie8 (ganeshie8):

Actually no, your method works perfectly fine for powers of 2 :)

ganeshie8 (ganeshie8):

i misinterpreted earlier :O

OpenStudy (mathmath333):

it works for 8 and 7

ganeshie8 (ganeshie8):

that looks interesting

OpenStudy (mathmath333):

yes i was just looking to quickly find the \(\huge a^{b^{c}}\) in case of \(\large (mod ~~10)\)

OpenStudy (mathmath333):

so i think misty's method was little dubious

ganeshie8 (ganeshie8):

i wouldnt doubt misty's method, all methods are giving the same answer... depends on which one you're more happy wid thats all :)

ganeshie8 (ganeshie8):

so you have a general strategy for finding the units digit of \(\large a^{b^c}\) for all \(a\) ? xD

OpenStudy (mathmath333):

yes but i m still confused why didnt it work for \(\Large 2^{{117}^{188}}\)

OpenStudy (mathmath333):

yes tactics rather than stratagy

ganeshie8 (ganeshie8):

looks like this mod4 strategy (tactic) works for all a ?

ganeshie8 (ganeshie8):

can we say \(\large a^{b^c} \) and \(\large a^{\frac{b^c}{4}}\) will have same units digit for all \(a\) ?

ganeshie8 (ganeshie8):

looks like it works !

ganeshie8 (ganeshie8):

it should work for a=3 also because 3^4 = 81 since the last digit is 1, it has to work

OpenStudy (mathmath333):

\(\LARGE \begin{align} \color{black}{ 7^{{152}^{188}} \hspace{.33em}\\~\\ \implies 7^{\frac{152^{188}}{4}} \hspace{.33em}\\~\\ \implies 7^{\frac{0^{188}}{4}} \hspace{.33em}\\~\\ \implies 7^{\frac{0}{4}} \hspace{.33em}\\~\\ \implies \dfrac{7^{4}}{10} \hspace{.33em}\\~\\ \implies \frac{1}{10} \hspace{.33em}\\~\\ \implies 1 \hspace{.33em}\\~\\ }\end{align}\) It will only work for \(2,7,8\) as they have a cycle of \(4\)

OpenStudy (mathmath333):

for three there is a more easy one !

ganeshie8 (ganeshie8):

it works for all \(a\) i agree we can have more easier methods for 3,4,5,6,9 :)

ganeshie8 (ganeshie8):

your method works for all \(a\)

ganeshie8 (ganeshie8):

show me a counter example if you can :)

OpenStudy (mathmath333):

well yes it works for 2,3,7,8 as they have a cycle of 4 but but in case of cycle of 2 there is another lazy tactic!

ganeshie8 (ganeshie8):

you want to take mod 2 for the exponent instead of mod 4 ?

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{ \normalsize \text{unit's digit of } \hspace{.33em}\\~\\ 9^{{152}^{188}} \hspace{.33em}\\~\\ 9^{odd}\implies 9 \hspace{.33em}\\~\\ 9^{even}\implies 1 \hspace{.33em}\\~\\ {152}^{188} \normalsize \text{its even } \hspace{.33em}\\~\\ 9^{{152}^{188}} \implies 9^{even}\implies 1 \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (mathmath333):

these will work for \(4~~ and ~~9\)

ganeshie8 (ganeshie8):

beautiful xD for 4 : 4^odd = 4 4^even = 6 is it ?

OpenStudy (mathmath333):

yes that one!

OpenStudy (mathmath333):

it saves time haha

ganeshie8 (ganeshie8):

Clever!

OpenStudy (mathmath333):

and no need for \(1,5~~and~~6\)